LOG-IN
Displaying reports 301-320 of 3360.Go to page Start 12 13 14 15 16 17 18 19 20 End
R&D (FilterCavity)
Print this report.
MichaelPage - 19:19, Tuesday 28 May 2024 (3572)Get code to link to this report
OPO recovery - initial Faraday insolator

Logan, Michael

Logan learned about how to adjust the transmission and blocking polarizer setup of the Faraday isolator. It seems everything was mostly in place regarding polarization - the required input is p-pol for IO-5-1064-HP. With current QWP/HWP setup about 300 to 1 p-pol vs s-pol could be reached, checked via PBS and power meter.

We thought there were some issues with the FI not transmitting close to 100% but this turned out to be putting the power meter too far upstream, and then also small alignment tweaks. I could get about 95%+ transmission just aligning by hand. I tried to tweak with the micro adjusters but it turns out the platform is stuck and adjusting yaw makes it bump upwards (hard to describe without a picture). We should buy a better mount.

The initial point of this exercise was to check the contribution of the FI to some weird possible HOM seen on the beam profiler, hence why I put the micro-adjuster underneath to adjust pitch and yaw of the FI. I think at this stage we should just get on with it - check the output of the FI with the beam profiler and see if we can get rid of the weird mode. Super optimal transmission is not so important.

General (General)
Print this report.
MarcEisenmann - 19:10, Tuesday 28 May 2024 (3571)Get code to link to this report
internet issues

[Hugo, Marc, Shalika]

Since sometimes now (maybe from last electric shutdown) one of the wifi relay in the central building broke (in storage) room.

In addition the ethernet switch seems broken as well. NOC is aware of this issue.

In the mean time, we placed a small switch on top of LC table and connected all TAMA pcs to it. We get internet from the new DGS switch.

Internet is back on all pc but at least PCI pc has now issue finding the Matlab license...

KAGRA MIR (Polarization)
Print this report.
MarcEisenmann - 19:08, Tuesday 28 May 2024 (3569)Get code to link to this report
birefringence measurement preparation

[Hugo, Marc, Shalika]

laser diode 1A : 1.6mW

z = 20 mm gives distance between blades and last mirror on injection = 171mm

 horizontal cut

Y: 103 mm cut (power halved) (power on dc 1.1) after hav power on dc was 0.55

X:374mm

z = 20 : Tue, May 28, 2024 2-14-49 PM.txt

z = 40 : Tue, May 28, 2024 2-21-29 PM.txt

z = 60: Tue, May 28, 2024 2-25-28 PM.txt

z = 80: Tue, May 28, 2024 2-29-48 PM.txt

 

verticale:

X: 313

Y : 191

z = 20 : Tue, May 28, 2024 11-32-57 AM.txt

z = 40 : Tue, May 28, 2024 2-01-31 PM.txt

z = 60: Tue, May 28, 2024 2-06-27 PM.txt

z = 80:Tue, May 28, 2024 2-10-41 PM.txt

after realignment horizontal

z = 20 : Tue, May 28, 2024 4-14-32 PM.txt

z = 40 : Tue, May 28, 2024 4-30-15 PM.txt

z = 60 : Tue, May 28, 2024 4-34-03 PM.txt

z = 80 : Tue, May 28, 2024 4-37-59 PM.txt

In the end we recovered normal incidence in both vertical and horizontal directions (within 0.026deg).

We wanted to prepare for linear polariscope measurement but the lockin amplifier showed strange error (undr) related to the chopper signal...

We'll start with polarization camera measurements in the coming days

R&D (FilterCavity)
Print this report.
MarcEisenmann - 19:02, Tuesday 28 May 2024 (3570)Get code to link to this report
Comment to DDS2 issues investigation (Click here to view original report: 3560)

[Logan, Marc]

We added one of the 2 remaining RF amplifier to the 500MHz clock output.

We soldered it to the NIM rack +12V and ground used for the LNVR. Note that we could have tried to use the +12V from the LNVR but were a bit worried to further damaged the clock..

The output of the clock is now +7.9dBm.

The output of the following splitter are

1 : -1.73dBm ; 2 : -6.19 dBm ; 3 : -5.09 dBm ; 4 : -5.33 dBm

Finally, we connected DDS2 to each output of the splitter and checked the signal output from DAC0.

We got about -9 dBm from all.

Next steps are fixing the clock crystal (was not fixed anymore) and the RF amplifier ; recover the electronic connections and recover squeezing.

BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 09:19, Monday 27 May 2024 (3567)Get code to link to this report
Comment to Optimization of Jones Matrix (Click here to view original report: 3564)

The birefringence properties remained same. Comes from the fact that we do J/Co

BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 17:27, Sunday 26 May 2024 (3566)Get code to link to this report
Comment to optimization of jones matrix (Click here to view original report: 3562)

after adding this, the jones matrix and info on birefringence was weird....I lost the crcuial info on the polarization properties rather than it working as a proper filter. So, not using this anymore.

Had actually picked it from a paper talking about improving noise in jones matrix. 

BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 17:23, Sunday 26 May 2024 (3565)Get code to link to this report
Comment to Optimization of Jones Matrix (Click here to view original report: 3564)

after this the jones matrix of mirror is not [1,0][0,-1] anymore. the 1's have changed to 0.8 ish. 

BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 08:24, Sunday 26 May 2024 (3564)Get code to link to this report
Optimization of Jones Matrix

Added power factor of sqrt(Output power/ Input Power) in jones vector. This is to take account the power getting halved after BS

Comments related to this report
ShalikaSingh - 17:23, Sunday 26 May 2024 (3565)

after this the jones matrix of mirror is not [1,0][0,-1] anymore. the 1's have changed to 0.8 ish. 

ShalikaSingh - 09:19, Monday 27 May 2024 (3567)

The birefringence properties remained same. Comes from the fact that we do J/Co

BIGFOOT (General)
Print this report.
HugoSaintOlive - 07:30, Saturday 25 May 2024 (3563)Get code to link to this report
Manipulation on the 2 AOM set up
I made the design of the set-up, I will use 4 different lenses to pass in the AOM probably, f=-50mm f=50mm f=-100mm f=200mm.


I removed the lens f=100mm and start to install the f=-50mm and f=50mm(not labeled yet) 10cm after the HWP.
BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 13:51, Friday 24 May 2024 (3562)Get code to link to this report
optimization of jones matrix

put the condition on jones vector such that matrix is computed only for 2 orthogonal set of jones vector. 

so the dot product is supposed to be 0. I have kept the range to be 85-90 deg. Therefore the dot product is required to be less than 0.2

Comments related to this report
ShalikaSingh - 17:27, Sunday 26 May 2024 (3566)

after adding this, the jones matrix and info on birefringence was weird....I lost the crcuial info on the polarization properties rather than it working as a proper filter. So, not using this anymore.

Had actually picked it from a paper talking about improving noise in jones matrix. 

BIGFOOT (General)
Print this report.
HugoSaintOlive - 10:16, Friday 24 May 2024 (3561)Get code to link to this report
New manipulation AOM
I use the begining of the set-up of the LCD and after the QWP and the HWP I put a lens f=100mm, and I characterize the beam.

Before the lens, w0 = -47,1 mm ; z0 = 26 μm , after the lens : w0 = 88.8 mm ; z0 = 49.2 μm

I moved the lens f=100mm after the characterization at d = 0.2 m (d = 0 is the same has the previous characterization)

!!!! The zero of reference did not correspond to any optic know, but it's at 101,5cm on the ruler, or the third a hole after the HWP.
R&D (FilterCavity)
Print this report.
MarcEisenmann - 20:43, Thursday 23 May 2024 (3560)Get code to link to this report
DDS2 issues investigation

[Logan, Marc]

Following the reinstallation of the mixers box we could not lock the GRMC.

We found out  it was because the DDS2 was not generating any sidebands but was just outputing some low amplitude broadband noise.

We could see that this issue was somehow related to the 500MHz clock as the software could not recognize it from time to time.

Today we checked the DDS2.

All voltages supply are correct.

All output (filtered or not) are just noise increase.

Then we checked the 500 MHz clock output and got

Channel 1 : -8.6dBm

Channel 2 : -16.7dBm

Channel 3 : -19dBm

Channel 4 : -21 dBm

We checked the output of Wenzel output and got -8dBm.

It seems there are some issues with the splitter (ZFRSC-4-842-S+) that creates the low clock power.

We swapped the clock of DDS2 to channel 1 and got the expected output powers.

We'll check if the remaining amplifier can be used for this clock output or if we have to purchase a new splitter.

Comments related to this report
MarcEisenmann - 19:02, Tuesday 28 May 2024 (3570)

[Logan, Marc]

We added one of the 2 remaining RF amplifier to the 500MHz clock output.

We soldered it to the NIM rack +12V and ground used for the LNVR. Note that we could have tried to use the +12V from the LNVR but were a bit worried to further damaged the clock..

The output of the clock is now +7.9dBm.

The output of the following splitter are

1 : -1.73dBm ; 2 : -6.19 dBm ; 3 : -5.09 dBm ; 4 : -5.33 dBm

Finally, we connected DDS2 to each output of the splitter and checked the signal output from DAC0.

We got about -9 dBm from all.

Next steps are fixing the clock crystal (was not fixed anymore) and the RF amplifier ; recover the electronic connections and recover squeezing.

R&D (FilterCavity)
Print this report.
LoganGarans - 14:51, Wednesday 22 May 2024 (3559)Get code to link to this report
Comment to OPO replacement ATC beam setup (Click here to view original report: 3556)
Because the beam profile was normal, I put back the QWP and the HWP and verified if it was s-polarized yesterday. Today, I put the new Faraday Isolator and adjusted it. I measured the intensity without the FI at around 30mW and with at around 27mW. That gives a transmission about 90% and the manufacturer gives 91% of transmission. But because I had trouble adjusting it, I'll do a beam profile tomorrow to check the new FI and the setup.
R&D (FilterCavity)
Print this report.
LoganGarans - 22:09, Tuesday 14 May 2024 (3558)Get code to link to this report
Comment to OPO replacement ATC beam setup (Click here to view original report: 3556)
Yesterday I prepared a new beam path to measure the beam profile directly from the laser. Today I measured the beam profile and at first it looks normal. So as we thought, the problem seems to be with the Faraday Isolator. I need to finish making the beam profile with all measurements to be sure of that but for now, it looks normal (at least with the W profile values).
R&D (FilterCavity)
Print this report.
MichaelPage - 18:45, Thursday 09 May 2024 (3557)Get code to link to this report
Comment to OPO replacement ATC beam setup (Click here to view original report: 3556)

We took a Thorlabs IO-5-1064-HP from a box of miscalleneous optics for KAGRA filter cavity testing (next to Mitsuhashi Roberts linkage experiment, purchased end of FY23).

We put it on a micro tilt adjustment stage and made sure it could be adjusted to the beam height from the Mephisto (72-73 mm). Actually the Faraday isolator that was already there was just put on a pedestal post instead of an adjustment stage, so it is likely that the vertical alignment through the FI was not good, generating a visible higher order mode and causing the improper beam profile.

R&D (FilterCavity)
Print this report.
MichaelPage - 16:23, Thursday 09 May 2024 (3556)Get code to link to this report
OPO replacement ATC beam setup

Logan, Michael

We once again found that the beam profile is showing some strange behavior with the asymmetry of the beam divergence. By using the profiler camera we see that it is probably due to misalignment of the Faraday isolator.

Comments related to this report
MichaelPage - 18:45, Thursday 09 May 2024 (3557)

We took a Thorlabs IO-5-1064-HP from a box of miscalleneous optics for KAGRA filter cavity testing (next to Mitsuhashi Roberts linkage experiment, purchased end of FY23).

We put it on a micro tilt adjustment stage and made sure it could be adjusted to the beam height from the Mephisto (72-73 mm). Actually the Faraday isolator that was already there was just put on a pedestal post instead of an adjustment stage, so it is likely that the vertical alignment through the FI was not good, generating a visible higher order mode and causing the improper beam profile.

LoganGarans - 22:09, Tuesday 14 May 2024 (3558)
Yesterday I prepared a new beam path to measure the beam profile directly from the laser. Today I measured the beam profile and at first it looks normal. So as we thought, the problem seems to be with the Faraday Isolator. I need to finish making the beam profile with all measurements to be sure of that but for now, it looks normal (at least with the W profile values).
LoganGarans - 14:51, Wednesday 22 May 2024 (3559)
Because the beam profile was normal, I put back the QWP and the HWP and verified if it was s-polarized yesterday. Today, I put the new Faraday Isolator and adjusted it. I measured the intensity without the FI at around 30mW and with at around 27mW. That gives a transmission about 90% and the manufacturer gives 91% of transmission. But because I had trouble adjusting it, I'll do a beam profile tomorrow to check the new FI and the setup.
BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 16:56, Tuesday 07 May 2024 (3553)Get code to link to this report
Comment to Compensation of BB1EO3 at 0 deg (Click here to view original report: 3544)

Did scans from

LC1, LC1step, LC2, LC2Step, averaging

0-25V, 0.2V, 0-25V, 0.2V

then observed and put a constraint on ell and azi of 2 deg

second scan 

0-1V, 0.01V, 0-25V, 0.1V

15-25V, 0.1V, 0-25V, 0.1V

third scan

0-0.06V, 0.01V, 0-8V, 0.05V

15-25V, 0.01V, 0-5.2V, 0.01V

15-25V, 0.01V, 5.2-8V, 0.05V

0-0.06V, 0.01V, 0-4.3V, 0.01V

4th scan

0-0.04V, 0.005V, 0.83-1.2V, 0.001V

15-25V, 0.005V, 0.83-1.2V, 0.001V

5th scan

14.99-15.02, 0.001, 0-2V, 0.001V

24-25V, 0.001V, 0.95-0.994V, 0.001V

BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 14:50, Sunday 05 May 2024 (3551)Get code to link to this report
Comment to Compensation of BB1EO3 at 0 deg (Click here to view original report: 3544)

Even though the voltages are being used from the file, it seems there could be a deviation, considering we did the calibration of LC 2 months ago. I can't see a zero crossing in azi and ell, also the file is a bit huge, so it will take longer than expected.

I am going to instead do scan using the voltage range estimated but without using the file, and doing first a rough scan, and then will later do a finer scan.

The voltage range estimated from characterization is 7-8 V for LC 1 and 0-25V for LC2 with condition of 2 degrees on azi and ell. 

So, doing a scan from 6-9V with 0.2V step, and 0-25V, 0,2V step. 

measurement file:

C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\AfterCompensation\Sun, May 5, 2024 2-44-03 PM.txt

(could see zero crossing in labview, still need to plot, measurement finished by the time I finished writing this elog)

BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 18:08, Saturday 04 May 2024 (3550)Get code to link to this report
Comment to Compensation of BB1EO3 at 0 deg (Click here to view original report: 3544)

Couldn't transfer voltage file , a bit too huge

decreased condition to 20mV, 2mV and 2 degrees.

Also, I had made a mistake in conversion from rad to degree in the previous one , at one instance

voltage file:

C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\CompensationVoltage\voltagecompensating_0.02_0.002_2.txt

measurement file :

C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\AfterCompensation\Sat, May 4, 2024 6-05-39 PM.txt

BIGFOOT (Cavity)
Print this report.
ShalikaSingh - 17:25, Saturday 04 May 2024 (3549)Get code to link to this report
Comment to Compensation of BB1EO3 at 0 deg (Click here to view original report: 3544)

The measurement stopped because of windows update. I have paused windows update for next 5 weeks.

Started another one with new voltage set of radius 50mV, step 5mV. The condition on ell and azi was 5 degrees.The measurment is being saved with 1 average.

voltage file is here:

C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\CompensationVoltage\voltagecompensating_0.05_0.005.txt

find the measurment file here:

C:\Users\atama\OneDrive\LC-Experiment\Measurement Data\Coating measurement\BB1-E03\20240417\0 deg\AfterCompensation\Sat, May 4, 2024 5-10-32 PM.txt